# Mathematical proof about spatial symmetry of solutions of the two-dimensional Kolmogorov flow

### v1

@@ Abstract
We give a mathematical proof that solution for all $t>0$ of the two-dimensional (2D) Kolmogorov flow governed by Navier-Stokes (NS) equations with periodic boundary condition remains the same spatial symmetry as its smooth initial condition at $t=0$. This mathematical theorem supports the corresponding CNS (clean numerical simulation) results of the 2D turbulent Kolmogorov flow [20, 8] that remain the same spatial symmetry, but does not support the corresponding DNS (direct numerical simulation) results that lose the spatial symmetry quickly. In other words, these DNS results violate this mathematical theorem. Thus, this mathematical theorem rigorously confirms that the spatiotemporal trajectories of NS turbulence given by DNS are indeed quickly polluted by numerical noises badly. It also illustrates that CNS can provide helpful enlightenments to deepen our understanding about turbulence and besides approach some mathematical truths about NS equations.
Keyword Navier-Stokes equations, spatial symmetry, turbulence, CNS, DNS
@@ 1 Introduction
Lorenz, who put forward in 1963 the famous “butterfly-effect” of chaos [14], further discovered in 2006 that numerical errors could have significant influences on global properties of chaotic systems [15], because unavoidable numerical noises as small artificial disturbance exponentially increase to macro-level. Note that the Navier-Stokes turbulence (i.e. turbulence governed by NS equations) is chaotic [5, 6, 2, 1, 3]. However, for all traditional numerical algorithms including the direct numerical simulation (DNS) of NS equations, numerical noises are unavoidable. Thus, logically speaking, numerical noises of DNS of NS turbulence should quickly increase to a macro-level, which certainly should lead to a distinct departure from exact trajectory of NS equations. Thus, DNS of NS turbulence should be quickly polluted by numerical noises badly.
@@
Note that the numerical noises of CNS are much smaller than those of DNS. So, CNS results can be used as benchmark solution to check the reliability of DNS results. In fact, some comparisons were given for two-dimensional (2D) incompressible Kolmogorov flow [20, 8]: the numerical simulations given by CNS remain the same spatial symmetry (in a finite but long enough time-interval) as the corresponding initial condition, but the corresponding DNS results remain the same spatial symmetry only at the beginning but quickly lose it. Thus, spatial symmetry becomes a key point to check the reliability of numerical simulation of NS turbulence.
In this paper, we give mathematical proofs that solution for all $t>0$ of the 2D Kolmogorov flow considered in [20, 8] remains the same spatial symmetry as its initial condition. Note that the CNS results in [20, 8] agree well with this proof, but the corresponding DNS results lose the spatial symmetry and thus violate this mathematical theorem, clearly confirming that DNS results are indeed quickly polluted by numerical noises badly. The detailed mathematical proofs are given in § 2, the concluding remarks and discussions are given in § 3.
2 Spatial symmetry of 2D Kolmogorov flow
@@
where $\psi$ is the stream-function, $t$ denotes the time, $x,y\in[0,2\pi]$ are horizontal and vertical coordinates, $Re$ is the Reynolds number, $n_{K}$ is an even number describing the Kolmogorov forcing scale, $\nabla^{2}$ is the Laplace operator, and $\nabla^{4}=\nabla^{2}\nabla^{2}$, respectively. The stream-function $\psi$ satisfies the periodic boundary condition
@@
Here, let us consider an arbitrarily smooth initial condition $\psi(x,y,0)$ that contains the spatial symmetry of rotation
@@
Assume that $\psi(x,y,t)$ can be expanded as Taylor series at $t=t_{0}\geq 0$ with a non-zero radius $\rho>0$ of convergence, say,
@@
where $|t-t_{0}|<\rho$, and
$\psi^{(m)}(x,y,t_{0})=\left.\frac{1}{m!}\frac{\partial^{m}\psi(x,y,t)}{\partial t^{m}}\right|_{t=t_{0}}$ (6)
with the definition
$\psi^{(0)}(x,y,t_{0})=\psi(x,y,t_{0}).$ (7)
Then, we can rewrite (1) as follows
@@
Under the coordinate transformation
@@
for the spatial symmetry of translation, respectively, the corresponding solution reads $\psi^{\prime}(x^{\prime},y^{\prime},t)$. Since there exists the spatial symmetry of rotation (3) and/or translation (4), it holds
@@
theorem 1
For the two-dimensional Kolmogorov flow governed by the Navier-Stokes equation (1) (with an even number $n_{K}$) under the periodic boundary condition (2) subject to a smooth initial condition with the spatial symmetry of rotation (3) and/or translation (4), if its solution $\psi(x,y,t)$ at $t=t_{0}\geq 0$ contains a spatial symmetry $\psi(x,y,t_{0})=\psi(x^{\prime},y^{\prime},t_{0})$ of rotation and/or translation, then there exists the spatial symmetry $\psi^{(m)}(x,y,t_{0})=\psi^{(m)}(x^{\prime},y^{\prime},t_{0})$ for all integers $m\geq 1$.
Proof: In the case of (9), we have
@@
$\frac{\partial f}{\partial x}\;\frac{\partial g}{\partial y}=\left(-\frac{\partial f}{\partial x^{\prime}}\right)\left(-\frac{\partial g}{\partial y^{\prime}}\right)=\frac{\partial f}{\partial x^{\prime}}\;\frac{\partial g}{\partial y^{\prime}}$ (16)
for smooth functions $f$ and $g$.
@@
$\frac{\partial f}{\partial x}\;\frac{\partial g}{\partial y}=\frac{\partial f}{\partial x^{\prime}}\;\frac{\partial g}{\partial y^{\prime}}$ (19)
for smooth functions $f$ and $g$.
Besides, in the case of (9), it holds
$\cos(n_{K}y^{\prime})=\cos(2n_{K}\pi-n_{k}y)=\cos(n_{K}y).$ (20)
In the case of (10), it holds
$\cos(n_{K}y^{\prime})=\cos(n_{K}\pi+n_{k}y)=\cos(n_{K}y),$ (21)
since $n_{k}$ is an even number.
Thus, in the case of (9) and/or (10), it holds
$\cos(n_{K}y)=\cos(n_{K}y^{\prime}),\;\;\nabla^{2}=\nabla^{\prime 2},\;\;\nabla^{4}=\nabla^{\prime 4},\;\;\frac{\partial f}{\partial x}\;\frac{\partial g}{\partial y}=\frac{\partial f}{\partial x^{\prime}}\;\frac{\partial g}{\partial y^{\prime}}$ (22)
for smooth functions $f$ and $g$.
@@
Since $\psi^{\prime}(x^{\prime},y^{\prime},t_{0})$ satisfies the same equations, we have
$\nabla^{\prime 2}\psi^{\prime(1)}(x^{\prime},y^{\prime},t)+\psi^{\prime(0)}_{x^{\prime}}\nabla^{\prime 2}\psi^{\prime(0)}_{y^{\prime}}-\psi^{\prime(0)}_{y^{\prime}}\nabla^{\prime 2}\psi^{\prime(0)}_{x^{\prime}}-\frac{1}{Re}\nabla^{\prime 4}\psi^{\prime(0)}+n_{K}\cos(n_{K}y^{\prime})=0.$ (24)
According to (22) and (23), we have at $t=t_{0}$ that
$\nabla^{\prime 2}\psi^{\prime(1)}(x^{\prime},y^{\prime},t)=\nabla^{2}\psi^{\prime(1)}(x^{\prime},y^{\prime},t),\;\;\;\cos(n_{K}y^{\prime})=\cos(n_{K}y)$ (25)
and
@@
$=$ $\psi^{\prime(0)}_{x^{\prime}}\left(\nabla^{\prime 2}\psi^{\prime(0)}\right)_{y^{\prime}}-\psi^{\prime(0)}_{y^{\prime}}\left(\nabla^{\prime 2}\psi^{\prime(0)}\right)_{x^{\prime}}-\frac{1}{Re}\nabla^{4}\psi^{\prime(0)}$
$=$ $\psi^{\prime(0)}_{x^{\prime}}\left(\nabla^{2}\psi^{\prime(0)}\right)_{y^{\prime}}-\psi^{\prime(0)}_{y^{\prime}}\left(\nabla^{2}\psi^{\prime(0)}\right)_{x^{\prime}}-\frac{1}{Re}\nabla^{4}\psi^{\prime(0)}$
@@
$\nabla^{2}\psi^{\prime(1)}(x^{\prime},y^{\prime},t)+\psi^{(0)}_{x}\nabla^{2}\psi^{(0)}_{y}-\psi^{(0)}_{y}\nabla^{2}\psi^{(0)}_{x}-\frac{1}{Re}\nabla^{4}\psi^{(0)}+n_{K}\cos(n_{K}y)=0.$ (27)
@@
$\psi^{\prime(1)}(x^{\prime},y^{\prime},t_{0})=\psi^{(1)}(x,y,t_{0}).$ (28)
(B) For a given positive integer $m$
Assume that for a given positive integer $m$ there exists the spatial symmetry at $t=t_{0}$, say,
$\psi^{\prime(i)}(x^{\prime},y^{\prime},t_{0})=\psi^{(i)}(x,y,t_{0}),\hskip 28.45274pt0\leq i\leq m.$ (29)
Differentiating (1) and (24) $m$ times with respect to $t$ and then dividing by $m!$, we have
$(m+1)\nabla^{2}\psi^{(m+1)}(x,y,t)$ (30)
$=$ $\frac{1}{Re}\nabla^{4}\psi^{(m)}(x,y,t)-\sum_{i=0}^{m}\Big[\psi^{(i)}_{x}\nabla^{2}\psi^{(m-i)}_{y}-\psi^{(i)}_{y}\nabla^{2}\psi^{(m-i)}_{x}\Big$]
and
$(m+1)\nabla^{\prime 2}\psi^{\prime(m+1)}(x^{\prime},y^{\prime},t)$ (31)
$=$ $\frac{1}{Re}\nabla^{\prime 4}\psi^{\prime(m)}(x^{\prime},y^{\prime},t)-\sum_{i=0}^{m}\left[\psi^{\prime(i)}_{x^{\prime}}\nabla^{\prime 2}\psi^{\prime(m-i)}_{y^{\prime}}-\psi^{\prime(i)}_{y^{\prime}}\nabla^{\prime 2}\psi^{\prime(m-i)}_{x^{\prime}}\right$],
respectively.
According to (22) and (29), we have at $t=t_{0}$ that
$\nabla^{\prime 2}\psi^{\prime(m+1)}=\nabla^{2}\psi^{\prime(m+1)},\;\;\;\nabla^{\prime 4}\psi^{\prime(m)}=\nabla^{4}\psi^{(m)},$ (32)
and
$\sum_{i=0}^{m}\left[\psi^{\prime(i)}_{x^{\prime}}\nabla^{\prime 2}\psi^{\prime(m-i)}_{y^{\prime}}-\psi^{\prime(i)}_{y^{\prime}}\nabla^{\prime 2}\psi^{\prime(m-i)}_{x^{\prime}}\right$] (33)
$=$ $\sum_{i=0}^{m}\left[\psi^{\prime(i)}_{x^{\prime}}\left(\nabla^{\prime 2}\psi^{\prime(m-i)}\right)_{y^{\prime}}-\psi^{\prime(i)}_{y^{\prime}}\left(\nabla^{\prime 2}\psi^{\prime(m-i)}\right)_{x^{\prime}}\right]$
$=$ $\sum_{i=0}^{m}\left[\psi^{\prime(i)}_{x^{\prime}}\left(\nabla^{2}\psi^{\prime(m-i)}\right)_{y^{\prime}}-\psi^{\prime(i)}_{y^{\prime}}\left(\nabla^{2}\psi^{\prime(m-i)}\right)_{x^{\prime}}\right]$
$=$ $\sum_{i=0}^{m}\left[\psi^{(i)}_{x}\left(\nabla^{2}\psi^{(m-i)}\right)_{y}-\psi^{(i)}_{y}\left(\nabla^{2}\psi^{(m-i)}\right)_{x}\right]$
$=$ $\sum_{i=0}^{m}\Big[\psi^{(i)}_{x}\nabla^{2}\psi^{(m-i)}_{y}-\psi^{(i)}_{y}\nabla^{2}\psi^{(m-i)}_{x}\Big$].
Substituting (32) and (33) into (31) gives at $t=t_{0}$ that
$(m+1)\nabla^{2}\psi^{\prime(m+1)}(x^{\prime},y^{\prime},t_{0})$ (34)
$=$ $\frac{1}{Re}\nabla^{4}\psi^{(m)}(x,y,t_{0})-\sum_{i=0}^{m}\Big[\psi^{(i)}_{x}\nabla^{2}\psi^{(m-i)}_{y}-\psi^{(i)}_{y}\nabla^{2}\psi^{(m-i)}_{x}\Big$].
Comparing (34) to (30) gives the spatial symmetry
$\psi^{\prime(m+1)}(x^{\prime},y^{\prime},t_{0})=\psi^{(m+1)}(x,y,t_{0})$ (35)
for a given positive integer $m$.
(C) For an arbitrary integer $m\geq 0$
According to (A) and (B), it holds the spatial symmetry
@@
for an arbitrary integer $m\geq 0$. This ends the proof of theorem 1.
theorem 2
For the two-dimensional Kolmogorov flow governed by the Navier-Stokes equation (1) (with an even number $n_{K}$) under the periodic boundary condition (2) subject to a smooth initial condition with the spatial symmetry of rotation (3) and/or translation (4), if its solution $\psi(x,y,t)$ at $t=t_{0}$ contains a spatial symmetry of rotation and/or translation, say, $\psi^{\prime}(x^{\prime},y^{\prime},t_{0})=\psi(x,y,t_{0})$, and besides the radius $\rho$ of convergence of its Taylor series at $t=t_{0}$ is non-zero, then $\psi(x,y,t)$ in $t\in[t_{0},t_{0}+\rho)$ has the same spatial symmetry as $\psi(x,y,t_{0})$.
Proof: According to theorem 1, from the spatial symmetry $\psi^{\prime}(x^{\prime},y^{\prime},t_{0})=\psi(x,y,t_{0})$, we have the spatial symmetry
@@
where $|t-t_{0}|<\rho$, including $t\in[t_{0},t_{0}+\rho)$, where $\rho>0$ is the radius of convergence. This guarantees the spatial symmetry $\psi^{\prime}(x^{\prime},y^{\prime},t)=\psi(x,y,t)$ in $t\in[t_{0},t_{0}+\rho)$. This ends the proof of theorem 2.
theorem 3
For the two-dimensional Kolmogorov flow governed by the Navier-Stokes equation (1) (with an even number $n_{K}$) under the periodic boundary condition (2) subject to a smooth initial condition with the spatial symmetry of rotation (3) and/or translation (4), if the convergence radius $\rho$ of its Taylor series is non-zero for all $t\geq 0$, its solution for all $t>0$ has the same spatial symmetry as its initial condition.
Proof: Since the initial condition contains the spatial symmetry, i.e. $\psi(x,y,0)=\psi^{\prime}(x^{\prime},y^{\prime},0)$, we have according to theorem 2 (by setting $t_{0}=0$) the spatial symmetry
@@
where $\delta_{0}<\rho_{0}$, and $\rho_{0}$ is the radius of convergence of its Taylor series at $t=0$. Then, setting $t_{0}=\delta_{0}$, we have according to theorem 2 the spatial symmetry
@@
where $\delta_{1}<\rho_{1}$, and $\rho_{1}$ is the radius of convergence of its Taylor series at $t=t_{0}=\delta_{0}$. Furthermore, setting $t_{0}=\delta_{1}$, we have according to theorem 2 the spatial symmetry
@@
$\psi(x,y,t)=\psi^{\prime}(x^{\prime},y^{\prime},t),\hskip 28.45274ptt\in[0,\delta_{2}].$
The same process can keep going, since the Taylor series of $\psi(x,y,t)$ always exists with a non-zero radius of convergence. Thus, it holds the spatial symmetry

### v2

# Spatial symmetry invariance of solution of Kolmogorov flow

@@ Abstract
We prove a mathematical theorem that solution for all $t>0$ of the two-dimensional (2D) Kolmogorov flow governed by Navier-Stokes (NS) equations with periodic boundary condition keeps the same spatial symmetry as its smooth initial condition. This mathematical theorem can be used to check the correctness and reliability of numerical simulations of NS turbulence. For example, it supports the corresponding CNS (clean numerical simulation) results of the 2D turbulent Kolmogorov flow [21, 8] that remain the same spatial symmetry in the whole time interval of simulation, but does not support the corresponding DNS (direct numerical simulation) results that lose the spatial symmetry quickly. In other words, these DNS results violate this mathematical theorem. Thus, this mathematical theorem rigorously confirms that the spatiotemporal trajectories of NS turbulence given by DNS are indeed quickly polluted by numerical noises badly. It also illustrates that CNS can provide helpful enlightenments to deepen our understanding about turbulence and besides approach some mathematical truths about NS equations.
Keyword Navier-Stokes equations, Kolmogorov flow, spatial symmetry, turbulence, CNS, DNS
MSC 76F02, 76F65, 65M25
@@ 1 Introduction
Lorenz, who put forward in 1963 the famous “butterfly-effect” of chaos [14], further discovered in 2006 that numerical errors could have significant influences on global properties of chaotic systems [15], because unavoidable numerical noises as small artificial disturbance exponentially increase to macro-level. Note that the Navier-Stokes (NS) turbulence (i.e. turbulence governed by NS equations) is chaotic [5, 6, 2, 1, 3]. However, for all traditional numerical algorithms including the direct numerical simulation (DNS) of NS equations, numerical noises are unavoidable. Thus, logically speaking, according to the butterfly-effect of chaotic system, numerical noises of DNS of NS turbulence should quickly increase to a macro-level, which certainly should lead to a distinct departure of numerical trajectory from exact solution of NS equations. Thus, logically speaking, DNS of NS turbulence should be quickly polluted by numerical noises badly.
@@
Note that the numerical noises of CNS are much smaller than those of DNS. So, CNS results can be used as benchmark solution to check the correctness and reliability of DNS results. In fact, some comparisons were given for two-dimensional (2D) incompressible Kolmogorov flow [21, 8]: the numerical simulations given by CNS remain the same spatial symmetry (in a finite but long enough time-interval) as the corresponding initial condition, but the corresponding DNS results remain the same spatial symmetry only at the beginning but quickly lose it. Thus, spatial symmetry becomes a key point to check the reliability of numerical simulation of NS turbulence.
In this paper, we prove a mathematical theorem that exact solution of the 2D Kolmogorov flow considered in [21, 8] always remains the same spatial symmetry for all $t>0$ as its initial condition. Note that the CNS results in [21, 8] agree well with this theorem, but the corresponding DNS results lose the spatial symmetry and thus violate this mathematical theorem, clearly confirming that DNS results are indeed quickly polluted by numerical noises badly, and depart distinctly away from exact solution. The detailed mathematical proofs are given in § 2, the concluding remarks and discussions are given in § 3.
2 Spatial symmetry invariance of solution
@@
where $\psi$ is the stream-function, $t$ denotes the time, $x,y\in[0,2\pi]$ are horizontal and vertical coordinates in the coordinate system OxyO-xy, ReRe is the Reynolds number, $n_{K}$ is an even number describing the Kolmogorov forcing scale, $\nabla^{2}$ is the Laplace operator, and $\nabla^{4}=\nabla^{2}\nabla^{2}$, respectively. The stream-function $\psi$ satisfies the periodic boundary condition
@@
Here, let us consider a smooth initial condition $\psi(x,y,0)$ that contains the spatial symmetry of rotation
@@
Define
$\psi^{(m)}(x,y,t)=\frac{1}{m!}\frac{\partial^{m}\psi(x,y,t)}{\partial t^{m}}$ (5)
with
$\psi^{(0)}(x,y,t)=\psi(x,y,t).$ (6)
Assume that $\psi(x,y,t)$ can be expanded as Taylor series at $t=t_{0}\in[0,+\infty)$ with a non-zero radius $\rho>0$ of convergence, say,
@@
where $|t-t_{0}|<\rho$. Then, we can rewrite (1) as follows
@@
Let us consider another coordinate system $O^{\prime}-x^{\prime}y^{\prime}$, where
@@
for translation, respectively, in which the corresponding solution reads $\psi^{\prime}(x^{\prime},y^{\prime},t)$. Since $\psi^{\prime}(x^{\prime},y^{\prime},t)$ satisfies the same equations, it holds
$\nabla^{\prime 2}\psi^{\prime(1)}+\psi^{\prime}_{x^{\prime}}\nabla^{\prime 2}\psi^{\prime}_{y^{\prime}}-\psi^{\prime}_{y^{\prime}}\nabla^{\prime 2}\psi^{\prime}_{x^{\prime}}-\frac{1}{Re}\nabla^{\prime 4}\psi^{\prime}+n_{K}\cos(n_{K}y^{\prime})=0$ (11)
under the periodic boundary condition, subject to an initial condition $\psi^{\prime}(x^{\prime},y^{\prime},0)$. Since there exists the spatial symmetry of rotation (3) and/or translation (4) for the initial solution, it holds
@@
Lemma 1
Let $n_{K}\geq 2$ denote an even integer, $f$ and $g$ denote two smooth functions with spatial symmetry $f(x,y,t)=f^{\prime}(x^{\prime},y^{\prime},t)$ and $g(x,y,t)=g^{\prime}(x^{\prime},y^{\prime},t)$, respectively. It holds
$\cos(n_{k}y)=\cos(n_{k}y^{\prime}),\;\;\;\nabla^{2}=\nabla^{\prime 2},\;\;\;\nabla^{4}=\nabla^{\prime 4},\;\;\;\frac{\partial f}{\partial x}\;\frac{\partial g}{\partial y}=\frac{\partial f^{\prime}}{\partial x^{\prime}}\;\frac{\partial g^{\prime}}{\partial y^{\prime}}$ (13)
in the case of (9) and/or (10).
Proof. In the case of (9), we have
$\cos(n_{K}y^{\prime})=\cos(2n_{K}\pi-n_{k}y)=\cos(n_{K}y),$ (14)
@@
$\frac{\partial f}{\partial x}\;\frac{\partial g}{\partial y}=\left(-\frac{\partial f^{\prime}}{\partial x^{\prime}}\right)\left(-\frac{\partial g^{\prime}}{\partial y^{\prime}}\right)=\frac{\partial f^{\prime}}{\partial x^{\prime}}\;\frac{\partial g^{\prime}}{\partial y^{\prime}}$ (19)
for two smooth functions $f,g$ with the spatial symmetry $f=f^{\prime}$ and $g=g^{\prime}$.
@@
$\frac{\partial f}{\partial x}\;\frac{\partial g}{\partial y}=\frac{\partial f^{\prime}}{\partial x^{\prime}}\;\frac{\partial g^{\prime}}{\partial y^{\prime}}$ (23)
for two smooth functions $f,g$ with the spatial symmetry $f=f^{\prime}$ and $g=g^{\prime}$. This ends the proof of Lemma 1.
Lemma 2
For the two-dimensional Kolmogorov flow governed by the Navier-Stokes equation (1) (with an even number $n_{K}$) under the periodic boundary condition (2) subject to a smooth initial condition with the spatial symmetry of rotation (3) and/or translation (4), if its solution $\psi(x,y,t)$ at $t=t_{0}\geq 0$ contains a spatial symmetry $\psi(x,y,t_{0})=\psi^{\prime}(x^{\prime},y^{\prime},t_{0})$ of rotation and/or translation, then there exists the spatial symmetry $\psi^{(m)}(x,y,t_{0})=\psi^{\prime(m)}(x^{\prime},y^{\prime},t_{0})$ for all integers $m\geq 1$.
Proof. This Lemma can be proved by means of the recursive method.
@@
According to (13) of Lemma 1 and (24), we have at $t=t_{0}$ that
@@
$=$ $\psi^{\prime(0)}_{x^{\prime}}\left(\nabla^{\prime 2}\psi^{\prime(0)}\right)_{y^{\prime}}-\psi^{\prime(0)}_{y^{\prime}}\left(\nabla^{\prime 2}\psi^{\prime(0)}\right)_{x^{\prime}}-\frac{1}{Re}\nabla^{4}\psi^{(0)}$
@@
$\nabla^{2}\psi^{\prime(1)}+\psi^{(0)}_{x}\nabla^{2}\psi^{(0)}_{y}-\psi^{(0)}_{y}\nabla^{2}\psi^{(0)}_{x}-\frac{1}{Re}\nabla^{4}\psi^{(0)}+n_{K}\cos(n_{K}y)=0.$ (26)
Comparing (26) to (8) gives the spatial symmetry at $t=t_{0}$ that
$\psi^{\prime(1)}(x^{\prime},y^{\prime},t_{0})=\psi^{(1)}(x,y,t_{0}).$ (27)
Thus, Lemma 2 holds true when $m=1$.
(B) Assume that Lemma 2 holds true when $m=1,2,\cdots,n$, say, there exists the spatial symmetry at $t=t_{0}$:
$\psi^{\prime(i)}(x^{\prime},y^{\prime},t_{0})=\psi^{(i)}(x,y,t_{0}),\hskip 28.45274pt0\leq i\leq n.$ (28)
Differentiating (1) and (11) $n$ times with respect to $t$ and then dividing by $n!$, we have
$(n+1)\nabla^{2}\psi^{(n+1)}=\frac{1}{Re}\nabla^{4}\psi^{(n)}-\sum_{i=0}^{n}\Big[\psi^{(i)}_{x}\nabla^{2}\psi^{(n-i)}_{y}-\psi^{(i)}_{y}\nabla^{2}\psi^{(n-i)}_{x}\Big$] (29)
in the coordinate system OxyO-xy, and
$(n+1)\nabla^{\prime 2}\psi^{\prime(n+1)}=\frac{1}{Re}\nabla^{\prime 4}\psi^{\prime(n)}-\sum_{i=0}^{n}\left[\psi^{\prime(i)}_{x^{\prime}}\nabla^{\prime 2}\psi^{\prime(n-i)}_{y^{\prime}}-\psi^{\prime(i)}_{y^{\prime}}\nabla^{\prime 2}\psi^{\prime(n-i)}_{x^{\prime}}\right$] (30)
in the coordinate system $O^{\prime}-x^{\prime}y^{\prime}$, respectively.
According to (13) of Lemma 1 and (28), we have at $t=t_{0}$ that
$\nabla^{\prime 2}\psi^{\prime(n+1)}=\nabla^{2}\psi^{\prime(n+1)},\;\;\;\nabla^{\prime 4}\psi^{\prime(n)}=\nabla^{4}\psi^{(n)},$ (31)
and
$\sum_{i=0}^{n}\left[\psi^{\prime(i)}_{x^{\prime}}\nabla^{\prime 2}\psi^{\prime(n-i)}_{y^{\prime}}-\psi^{\prime(i)}_{y^{\prime}}\nabla^{\prime 2}\psi^{\prime(n-i)}_{x^{\prime}}\right$] (32)
$=$ $\sum_{i=0}^{n}\left[\psi^{\prime(i)}_{x^{\prime}}\left(\nabla^{\prime 2}\psi^{\prime(n-i)}\right)_{y^{\prime}}-\psi^{\prime(i)}_{y^{\prime}}\left(\nabla^{\prime 2}\psi^{\prime(n-i)}\right)_{x^{\prime}}\right]$
$=$ $\sum_{i=0}^{n}\left[\psi^{(i)}_{x}\left(\nabla^{2}\psi^{(n-i)}\right)_{y}-\psi^{(i)}_{y}\left(\nabla^{2}\psi^{(n-i)}\right)_{x}\right]$
$=$ $\sum_{i=0}^{n}\Big[\psi^{(i)}_{x}\nabla^{2}\psi^{(n-i)}_{y}-\psi^{(i)}_{y}\nabla^{2}\psi^{(n-i)}_{x}\Big$].
Substituting (31) and (32) into (30) gives
$(n+1)\nabla^{2}\psi^{\prime(n+1)}=\frac{1}{Re}\nabla^{4}\psi^{(n)}-\sum_{i=0}^{n}\Big[\psi^{(i)}_{x}\nabla^{2}\psi^{(n-i)}_{y}-\psi^{(i)}_{y}\nabla^{2}\psi^{(n-i)}_{x}\Big$]. (33)
@@
$\psi^{\prime(n+1)}(x^{\prime},y^{\prime},t_{0})=\psi^{(n+1)}(x,y,t_{0})$ (34)
for integer $n\geq 1$. Thus, if Lemma 2 holds for an integer $n\geq 1$, it also holds true for the integer $n+1$.
(C) Based on (A) and (B), we have due to the the recursive method that the spatial symmetry
@@
holds for an arbitrary integer $m\geq 1$. This ends the proof of Lemma 2.
Lemma 3
For the two-dimensional Kolmogorov flow governed by the Navier-Stokes equation (1) (with an even number $n_{K}$) under the periodic boundary condition (2) subject to a smooth initial condition with the spatial symmetry of rotation (3) and/or translation (4), if its solution $\psi(x,y,t)$ contains a spatial symmetry at $t=t_{0}$, say, $\psi^{\prime}(x^{\prime},y^{\prime},t_{0})=\psi(x,y,t_{0})$, and besides its Taylor series exists at $t=t_{0}$ with a non-zero radius $\rho$ of convergence, then $\psi(x,y,t)$ in $t\in[t_{0},t_{0}+\rho)$ has the same spatial symmetry as $\psi(x,y,t_{0})$.
Proof: According to Lemma 2, from the spatial symmetry $\psi^{\prime}(x^{\prime},y^{\prime},t_{0})=\psi(x,y,t_{0})$, we have the spatial symmetry
@@
where $|t-t_{0}|<\rho$, including $t\in[t_{0},t_{0}+\rho)$, where $\rho>0$ is the radius of convergence. This guarantees the spatial symmetry $\psi^{\prime}(x^{\prime},y^{\prime},t)=\psi(x,y,t)$ in $t\in[t_{0},t_{0}+\rho)$. This ends the proof of Lemma 3.
Theorem of spatial symmetry invariance
For the 2D Kolmogorov flow governed by the Navier-Stokes equation (1) (with an even number $n_{K}$) under the periodic boundary condition (2) subject to a smooth initial condition with the spatial symmetry of rotation (3) and/or translation (4), if its Taylor series exists with a non-zero convergence radius $\rho$ for all $t\geq 0$, its solution for all $t>0$ keeps the same spatial symmetry as its initial condition.
Proof: Since the initial condition contains the spatial symmetry, i.e. $\psi(x,y,0)=\psi^{\prime}(x^{\prime},y^{\prime},0)$, we have according to Lemma 3 (by setting $t_{0}=0$) the spatial symmetry
@@
where $\delta_{0}<\rho_{0}$, and $\rho_{0}$ is the radius of convergence of its Taylor series at $t=0$. Then, setting $t_{0}=\delta_{0}$, we have according to Lemma 3 the spatial symmetry
@@
where $\delta_{1}<\rho_{1}$, and $\rho_{1}$ is the radius of convergence of its Taylor series at $t=t_{0}=\delta_{0}$. Furthermore, setting $t_{0}=\delta_{1}$, we have according to Lemma 3 the spatial symmetry
@@
$\psi(x,y,t)=\psi^{\prime}(x^{\prime},y^{\prime},t),\hskip 28.45274ptt\in[0,\delta_{2}]$
in a larger interval of time. The same process can keep going, since the Taylor series of $\psi(x,y,t)$ always exists with a non-zero radius of convergence. Thus, it always holds the spatial symmetry